if x=acost y=asint, then b^2 x^2+a^2 y^2=
a)ab b)b^2+a^2
c)a^2b^2 d)a^4b^4

Grade:9

1 Answers

sonika p
6 years ago
x=acost
x^2 = a^2 cos^2 t
y = bsint
y^2 = b^2 sin^2 t
b^2 x^2+a^2 y^2=b^2*a^2cos^2 t +b^2*a^2sin^2 t = b^2*a^2[cos^2 t + sin^2 t] = b^2*a^2
as cos^2 t + sin^2 t =1
therefore, answer is option C